OpenStudy (callisto):

What's wrong with this? 4 = 4 [tan(π/4)] = 4sin(π/4) / cos(π/4) = 2 sin(π/2) / cos^2 (π/4) = sin(π) / [cos^2 (π/4) cos(π/2)] = 0/ 0?!

OpenStudy (blockcolder):

cos(pi/2) is zero, so you multiplied something by 0/0. =))

OpenStudy (callisto):

But the point is how to correct it?

OpenStudy (anonymous):

For this step : sin(π) / [cos^2 (π/4) cos(π/2)] we should not multiply 0/0 4sin(π/4) / cos(π/4) =2sin(π/2)/cos^2 (π/4) =2sin(π/2)/ (1/2)(cos(π/2) +1) =4sin(π/2) =4

OpenStudy (callisto):

I see.... So the double angle formula has its 'limitation' ... I'll not make the same mistake again! Thank you :)

OpenStudy (anonymous):

4 = 4 tan(π/4) 1 = tan (π/4) tan (π/4) = tan(π/4)

OpenStudy (callisto):

That's why I chose tan(π/4) :P

OpenStudy (anonymous):

it should be limitation of fraction :D

OpenStudy (callisto):

lol My math is not good. Sorry... limitation of fraction!~

OpenStudy (anonymous):

Oops, it's so obvious! tan (π/4) = 1

OpenStudy (anonymous):

Hey, think simple, will you! 1 = tan(π/4) = sin(π/4)/ cos(π/4) = 1 always true :P

OpenStudy (callisto):

I know~ But I was getting 4 = 0/0 :S Now, I know what's wrong :)